the ages of sarah's mother and aunt total 83 years. if her mother's age were doubled, then her mother would be 58 years older than her aunt. how old are sarah's mother and aunt.
Answer by JulietG(1812)

M + A = 83
2M = A + 58
therefore, 2M - 58 = A
substitute the value of A from the second equation into the first
M + (2M-58) = 83
3M - 58 = 83
3M = 141
M = 47
If her mother is 47, then her aunt is 83-47 = 36